### v2

@@ Abstract
Generating landscape paintings expands the possibilities of artistic creativity and imagination. Traditional landscape painting methods involve using ink or colored ink on rice paper, which requires substantial time and effort. These methods are susceptible to errors and inconsistencies and lack precise control over lines and colors. This paper presents LPGen, a high-fidelity, controllable model for landscape painting generation, introducing a novel multi-modal framework that integrates image prompts into the diffusion model. We extract its edges and contours by computing canny edges from the target landscape image. These, along with natural language text prompts and drawing style references, are fed into the latent diffusion model as conditions. We implement a decoupled cross-attention strategy to ensure compatibility between image and text prompts, facilitating multi-modal image generation. A decoder generates the final image. Quantitative and qualitative analyses demonstrate that our method outperforms existing approaches in landscape painting generation and exceeds the current state-of-the-art. The LPGen network effectively controls the composition and color of landscape paintings, generates more accurate images, and supports further research in deep learning-based landscape painting generation.
@@
image generation; decoupled cross-attention; latent diffusion model; controllability;
@@ 1 Introduction
Landscape painting Andrews (1999); Clunas (1997); Conisbee (1996); Way et al. (2001) holds a significant place in traditional art, depicting natural scenes to convey the artist’s aesthetic vision and emotional connection to nature. This genre utilizes tools such as ink, water, paper, and brushes, emphasizing artistic conception, brush and ink intensity variations, balanced composition, and effective use of negative space. With advances in computer technology, the generation of landscape paintings through modern algorithms has become a promising area of research.
@@
The creative process of landscape painting Guo (1999); Sze (1992) involves several intricate stages: outlining, chapping, rubbing, moss-dotting, and coloring, as shown in Figure 1 (a). Each stage requires a high level of skill and experience. Outlining requires precise control of lines, chapping captures the diverse textures of nature, rubbing, and dyeing necessitate meticulous control over the ink’s intensity and moisture, moss-dotting calls for a flexible and natural technique, and coloring requires a keen sense of color harmony. The complexity and precision involved in each stage contribute to the overall challenge of creating landscape paintings. Recent technological advancements, particularly in deep learning, have empowered computers to emulate landscape paintings’ artistic styles and brushwork. These technologies can generate artworks that capture the rich cultural essence of China, preserving traditional art forms and opening new paths for creative expression. This technological innovation supports the preservation and evolution of conventional culture Chen (2012); Williams and Marsh (2014). Research into the generation of landscape paintings is primarily divided into traditional and deep learning-based methods. Traditional methods include non-photorealistic rendering, image-based approaches, and computer-generated simulations. Non-photorealistic rendering Way et al. (2006, 2001); Liu (2015) uses computer graphics to simulate various artistic effects, such as ink diffusion, through specialized algorithms. Image-based methods Elgammal et al. (2017); Wang et al. (2018); Gatys et al. (2016) involve collecting brush stroke texture primitives (BSTP) Yu et al. (2003) from hand-drawn samples and employing them to map multiple layers to create mountain imagery. Despite their sophistication, image-based and non-photorealistic rendering methods are inherently complex, and their analog numerical modeling quality often falls short. Currently, many researchers are using generative adversarial networks (GANs) Goodfellow et al. (2014) to generate landscape paintings. However, due to the distinctive artistic style of landscape painting, GAN-based methods often encounter challenges such as unclear outlines, uncontrollable details, and inconsistent style migration.
{adjustwidth}
[…]
@@
Figure 1: The creative process of landscape painting involves several key steps: outlining, chapping, rubbing, moss-dotting, and coloring. Utilizing the stable diffusion method, landscape paintings can be precisely and flexibly generated by specifying a style reference image, a canny outline image, and appropriate cue words.
@@
By integrating computer vision with generative modeling techniques, such as stable diffusion (SD) Rombach et al. (2022), the artistic effects of landscape painting can be effectively simulated and reproduced Li et al. (2023); Sun et al. (2024). Diffusion models (DMs) have been successfully applied across various domains, including clothing swapping, face generation, dancing, and anime creation, demonstrating their versatility in generating high-quality realistic content from diverse inputs. For instance, diffusion models can produce highly realistic images nearly indistinguishable from actual photographs in face generation. Similarly, in anime creation, they can generate detailed and stylistically consistent images from simple prompts Lee et al. (2023); Kim et al. (2023); Choi et al. (2023); Yang et al. (2023). This method not only preserves the stylistic features of traditional landscape paintings but also allows precise control over the content and layout. This technology has broad applications in the creation of art, the preservation of cultural heritage, and digital display, as illustrated in Figure  1(b).
@@
Composition and color control are crucial factors in landscape painting, significantly influencing the overall effect and quality of the artwork. Previous style migration methods have often yielded unsatisfactory results in generating landscape paintings. To address this, we propose the LPGen framework, which draws inspiration from the fundamental steps of traditional landscape painting. This framework employs the canny edge detection algorithm to generate line drawings as input, a structure controller module to guide the image generation based on the canny outlines, and a style controller module to manage the painting style of the generated image. The main contributions of this work are summarized as follows:
@@
    We construct a comprehensive dataset comprising 2,520 high-resolution images, meticulously categorized into four distinct styles: azure green landscape, golden splendor landscape, ink wash landscape, and light vermilion landscape.
@@ 2.1 GAN-Based Method
Due to the rapid advancements in deep learning, significant progress has been made in various visual tasks Shen et al. (2023a, b). Based on the input content, landscape painting generation methods can be categorized into three types: nothing-to-image, image-to-image, and text-to-image.
For the first type, it does not require any input information to generate landscape paintings. An example is the Sketch-And-Paint GAN (SAPGAN) Xue (2021), the first neural network model capable of automatically generating traditional landscape paintings from start to finish. SAPGAN consists of two GANs: SketchGAN for generating edge maps and PaintGAN for converting these edge maps into paintings. Another example is an automated creation system Luo et al. (2022) based on GANs for landscape paintings, which comprises three cascading modules: generation, scaling, and super-resolution.
For the second type, a photo is input for style transfer. An interactive generation method Zhou et al. (2019) allows users to create landscape paintings by simply outlining with essential lines, which are then processed through a recurring adversarial generative network (RGAN) model to generate the final image. Another approach, neural abstraction style transfer Li et al. (2019), leverages the MXDoG filter and three fully differentiable loss terms. The ChipGAN architecture He et al. (2018) is an end-to-end GAN model that addresses critical techniques in ink painting, such as blanks, brush strokes, ink tone and spread. Additionally, an image style transfer framework Peng et al. (2022) utilizing CycleGAN, facilitates the conversion of photographs into paintings. The bilateral generative adversarial network algorithm (BiTGAN) He et al. (2023) constructs a bilateral generator using a shared encoder, U-Net path, ResNet path and decoder, incorporating an attentional network Shen et al. (2023) to convert photos into paintings. Furthermore, an attentional wavelet network Wang et al. (2021) utilizes wavelets to capture high-level mood and local details of paintings via the 2-D Haar wavelet transform. Lastly, the Paint-CUT model Wang et al. (2023) intelligently creates landscape paintings by utilizing a shuffle attentional residual block and edge enhancement techniques.
For the third type, the output image is guided by textual input information. A novel system Yuan et al. (2023) for Polaca transforms classical poetry into corresponding artistic landscape paintings and calligraphic works. Another approach, controlled landscape painting generation (CCLAP) Wang et al. (2023), is based on the latent diffusion model (LDM) Ho et al. (2020) and comprises two sequential modules: a content generator and a style aggregator. Although this method allows text-based control of the image, it still faces challenges related to the randomness of image generation and limited controllability.
@@ 2.2 Diffusion-Based Method
DM Ho et al. (2020) is a deep learning framework primarily utilized for image processing and computer vision tasks. Convert an input image into Gaussian noise through diffusion steps, and then reconstruct the original image via an inverse process. DM can generate detailed images from textual descriptions and are also effective for image restoration, image drawing, text-to-image, and image-to-image tasks Chen et al. (2024); Lian et al. (2024); Shen et al. (2023, 2024). Essentially, by providing a textual description of the desired image, SD can produce a realistic image that aligns with the given description Ho et al. (2020). These models reframe the "image generation" process into a "diffusion" process that incrementally removes noise. Starting with random Gaussian noise, the models progressively eliminate it through training until it is entirely removed, yielding an image that closely matches the text description Ho et al. (2020); Lian et al. (2024). However, a significant drawback of this approach is the considerable time and memory required, especially for high-resolution image generation Chen et al. (2024). LDM Rombach et al. (2022) was developed to address these limitations by significantly reducing memory and computational costs. This is achieved by applying the diffusion process within a lower-dimensional latent space rather than the high-dimensional pixel space Chen et al. (2024).
The emergence of large text-to-image models that can create visually appealing images from brief descriptive prompts has highlighted the remarkable potential of artificial intelligence. However, these models encounter challenges such as uneven data availability across specific domains compared to the generalized image-to-text domain. In addition, some tasks require more precise control and guidance than simple prompts can provide. ControlNet Zhang et al. (2023) addresses these issues by generating high-quality images based on user-provided cues and controls, which can fine-tune performance for specific tasks. Similarly, Ip-adapter Ye et al. (2023) employs decoupled cross-attention mechanisms for the characteristics of the text and the image. Inspired by these approaches, our study incorporates ControlNet’s additional structural controls, such as sketches, depth maps, and semantic segmentation maps. Moreover, the Ip-adapter can utilize style or content from reference images to guide the generation of landscape paintings. For example, the Thangka sketch colorization method Li et al. (2023) proposed by Li et al. employs multi-level adaptive instance-normalized color fusion combined with skip connection attention to effectively capture and distinguish the intricate semantic information embedded within the sketch lines. Additionally, Edge-MSGAN Sun et al. (2024) harnesses edge images to guide a completion network in generating complete ancient paintings. It employs a multibranch color correction network to precisely adjust and harmonize the colors, ensuring fidelity to the original artistic style.
@@ 3 Proposed Method
To address the above issues, LPGen has been developed to generate landscape paintings from photos. This section provides a comprehensive introduction to LPGen, detailing an overview of the model, the image structure controller, which manages the composition of the generated image, and the style controller, which dictates the painting style of the resulting image.
{adjustwidth}
[…]
Figure 2: Our proposed LPGen framework for landscape painting generation comprises three key components. First, the image structure controller module utilizes the Decoupled Cross-Attention technique to separately manage the structural and stylistic information of an image across different domains, allowing for precise control and regulation of specific elements and attributes in the generated image. Second, the style controller dynamically adjusts the features of the input image, enabling the generative model to accurately capture and reflect the style and structure of the source image. Finally, the text expression controller guides the image generation model using natural language descriptions, facilitating the production of high-quality images that align with the provided textual prompts.
@@ 3.1 Preliminaries
The complete structure of the LPGen framework is illustrated in Figure 2. The framework comprises three main components: the stable diffusion model, the structure controller, and the style controller. While the stable diffusion model offers generalized image generation capabilities, it lacks precise control over the structural and stylistic features of an image. To address this, we employ the structure controller to ensure that edge information effectively guides the image generation process. This controller guarantees that the generated image aligns with the composition of the canny reference image. Artists can dictate the position and shape of graphic elements through outlining, and LPGen uses the structure controller to manage the structure and form of the resulting image. For stylistic aspects, the style controller extracts the style of a reference segmented image and applies it to the generated image, allowing the generation of images in a specified style. The style controller learns various color schemes to convey different emotions. Building upon the stable diffusion model’s base image generation capabilities, the structure controller manages the structural layout, and the style controller governs the color scheme features. This precise and robust control enables us to achieve excellent management over the style and quality of the generated images, surpassing the capabilities of simple text-to-image models.
@@ 3.2 Structure Controller
The line drawing serves as the fundamental framework of landscape painting, establishing the overall layout and positioning of elements. During the coloring stage, the artist sequentially fills and renders the picture, coloring each element according to the guidelines set by the line drawing. To enable the line drawing to guide LDM in generating the image, the structure controller manipulates the neural network structure of the diffusion model by incorporating additional conditions. This, in combination with Stable Diffusion, ensures accurate spatial control, effectively addressing the spatial consistency issue. The image generation process thus emulates the actual painting process, transitioning from line drawing to coloring. We employ the structure controller as a model capable of adding spatial control to a pre-trained diffusion model beyond basic textual prompts. This controller integrates the UNet architecture from Stable Diffusion with a trainable UNet replica. This replica includes zero convolution layers within the encoder and middle blocks. The complete process executed by the structure controller is as follows:
@@ 3.3 Style Controller
In the original SD base model with text-generated graph capability, we introduce the style controller structure to integrate text cueing features, structure cueing features and image style cueing features. To achieve the above purpose, our core design is a decoupled cross-attention mechanism as shown in Figure 2 , where image features are embedded through a newly added cross-attention layer, which consists mainly of two modules: an image encoder for extracting image features, and an adaptation module with decoupled cross-attention for embedding the image features into a pre-trained text-to-image diffusion model.
The pre-trained CLIP image encoder model is used, but here, in order to efficiently decompose the global image Embedding, we use a small trainable projection network to project the image Embedding into a sequence of features of length N=4, which is the Image Feature in Fig. The projection network here is designed as a linear layer Linear plus a normalization layer LN, and at the same time, the dimensions of the input image features are kept consistent with the dimensions of the text features in the pre-trained diffusion model.
@@
$\mathcal{Z}_{new}=\mathcal{\textit{Attention}}(Q,K^{t},V^{t})+\lambda\cdot\mathcal{\textit{Attention}}(Q,K^{i},V^{i}),$ (2)

### v3

@@ Abstract
The creation of landscape paintings is a meticulous and skill-intensive endeavor, demanding exceptional artistic proficiency. With the development of artificial intelligence (AI) diffusion models, using these models for generating landscape paintings has become a novel method for addressing the aforementioned challenges. Herein, we build a unique traditional landscape painting dataset. Besides, we propose a high-fidelity, controllable model for landscape painting generation (LPGen), pioneering a novel multi-modal framework by incorporating the structure controller and the style controller into the diffusion model. Specifically, the structure controller module utilizes the Decoupled Cross-Attention technique to separately manage the structural information of an image across different domains, allowing for precise control and regulation of specific elements and attributes in the generated image. Meanwhile, the style controller dynamically adjusts the features of the input image, enabling the generative model to accurately capture and reflect the style of the canny image. The research results indicate that the proposed method outperforms existing approaches in landscape painting generation and exceeds the current state-of-the-art. Additionally, the proposed LPGen network excels in controlling the structure and style of landscape paintings, producing highly accurate images, and fostering further research in deep learning-based landscape painting generation.
@@
image generation; decoupled cross-attention; latent diffusion model; controllability
@@ 1 Introduction
Landscape painting holds a significant place in traditional art, depicting natural scenes to convey the artist’s aesthetic vision and emotional connection to nature Yao et al. (2024); Maggos et al. (2019); Chakhim et al. (2020); Asperti et al. (2023); Wang et al. (2023); Zengguo Sun (2024). This genre utilizes tools such as ink, water, paper, and brushes, emphasizing artistic conception, brush and ink intensity variations, balanced composition, and effective use of negative space Zengguo Sun (2024); Teresa Kurkiewicz (2023). With the development of computer technology, especially diffusion, the generation of landscape paintings through modern algorithms has become a promising area of research Yao et al. (2024); Lin et al. (2018); Sun et al. (2024).
The creative process of landscape painting involves several intricate stages: outlining, chapping, rubbing, moss-dotting, and coloring, as shown in Figure 1 (a) Zengguo Sun (2024). Outlining is the initial stage where the artist sketches the basic structure and composition of the landscape Zengguo Sun (2024); Yao et al. (2024); Teresa Kurkiewicz (2023). Chapping involves the careful depiction of textures, requiring an understanding of how to use brushstrokes to convey different natural surfaces Teresa Kurkiewicz (2023). Rubbing and dyeing necessitate meticulous control over ink intensity and moisture, balancing between depth and subtlety in shading Zengguo Sun (2024). Moss-dotting requires a flexible and natural technique to create the illusion of texture and dimension, adding complexity and life to the painting Teresa Kurkiewicz (2023). Finally, coloring demands a keen sense of color harmony to ensure the overall unity and aesthetic appeal of the artwork Yao et al. (2024). The complexity and precision involved in each stage contribute to the overall challenge of creating landscape paintings Zengguo Sun (2024); Yao et al. (2024); Teresa Kurkiewicz (2023).
@@
Recent technological advancements, particularly in deep learning, have empowered computers to emulate landscape paintings’ artistic styles and brushwork. These technologies can generate artworks that capture the rich cultural essence of China, preserving traditional art and opening new paths for creative expression Yang and Hu (2022); Li (2024); Way et al. (2023); Sun et al. (2023). This technological innovation supports the preservation and evolution of conventional culture Way et al. (2023); Sun et al. (2023); Way et al. (2023). Research into the generation of landscape paintings is primarily divided into traditional and deep learning-based methods. Traditional methods include non-photorealistic rendering, image-based approaches, and computer-generated simulations. Non-photorealistic rendering uses computer graphics to simulate various artistic effects, such as ink diffusion, through specialized algorithms Xin Zhang (2022); Andreas Semmo (2022). Image-based methods involve collecting brush stroke texture primitives (BSTP) from hand-drawn samples and employing them to map multiple layers to create mountain imagery Junho Kim (2024); Lu and Finkelstein (2023). Despite their sophistication, the quality of computer-generated simulations often falls short Xin Zhang (2022); Grubert et al. (2023).
@@
Figure 1: Hand-drawing method compared with the proposed LPGen method. The hand-drawing method involves a complex and repetitive process with several key steps such as outlining, chapping, rubbing, moss-dotting, and coloring. LPGen simplifies the process, making it convenient and flexible by precisely generating landscape paintings from a specified style reference image and a canny outline image.
@@
Currently, many researchers are using generative adversarial networks (GANs) to generate landscape paintings Park et al. (2021); Tang et al. (2022); Lee and Kang (2022); Zhang et al. (2023). However, due to the distinctive artistic style of landscape painting, GAN-based methods often face challenges such as unclear outlines, uncontrollable details, and inconsistent style migration Yang and Huang (2022). By visual generation, such as stable diffusion (SD), the artistic effects of landscape painting can be effectively simulated and reproduced Rombach et al. (2022); Liang and Zhao (2023). Diffusion models (DMs) have been successfully applied across various domains, including clothing swapping, face generation, dancing, and anime creation, demonstrating their versatility in generating high-quality realistic content from diverse inputs Ho et al. (2020); Meng and Zhang (2023); Liu and Chen (2023). For instance, diffusion models can produce highly realistic images nearly indistinguishable from actual photographs in face generation Choi and Lee (2023); Yang and Liu (2023). Similarly, in anime creation, they can generate detailed and stylistically consistent images from simple prompts Liu and Chen (2023); Park and Kim (2023). Although this method preserves the stylistic features of traditional landscape paintings, it does not allow precise control over the structure and style Liang and Zhao (2023); Rombach et al. (2022).
@@
Structure and style control are crucial factors in landscape painting, significantly influencing the overall effect and quality of the artwork. Previous GAN-based and Diffusion-Based methods have often yielded unsatisfactory results in generating landscape paintings. To address this, we propose the LPGen framework, which draws inspiration from the fundamental steps of traditional landscape painting. The framework consists of two parts: structure controller and style controller. The framework uses canny line graphs to control the painting structure through the structure controller and uses the reference style graph to control the style of the generated painting through the style controller. The main contributions of this work are summarized as follows:
@@
    We construct a comprehensive dataset comprising 2,416 high-resolution images, meticulously categorized into three distinct styles: azure green landscape, ink wash landscape, and light vermilion landscape.
@@ 2.1 GAN-Based Method
Due to the rapid advancements in deep learning, significant progress has been made in various visual tasks Shen et al. (2023a, b). Based on the input content, landscape painting generation methods can be categorized into three types: nothing-to-image, image-to-image, and text-to-image. For the first type, it does not require any input information to generate landscape paintings. An example is the Sketch-And-Paint GAN (SAPGAN) Xue (2021), the first neural network model capable of automatically generating traditional landscape paintings from start to finish. SAPGAN consists of two GANs: SketchGAN for generating edge maps and PaintGAN for converting these edge maps into paintings. Another example is an automated creation system Luo et al. (2022) based on GANs for landscape paintings, which comprises three cascading modules: generation, scaling, and super-resolution.
For the second type, a photo is input for style transfer. An interactive generation method Zhou et al. (2019) allows users to create landscape paintings by simply outlining with essential lines, which are then processed through a recurring adversarial generative network (RGAN) model to generate the final image. Another approach, neural abstraction style transfer Li et al. (2019), leverages the MXDoG filter and three fully differentiable loss terms. The ChipGAN architecture Sun et al. (2022) is an end-to-end GAN model that addresses critical techniques in ink painting, such as blanks, brush strokes, ink tone and spread. Furthermore, an attentional wavelet network Sun et al. (2024) utilizes wavelets to capture high-level mood and local details of paintings via the 2-D Haar wavelet transform. Lastly, the Paint-CUT model Wang et al. (2023) intelligently creates landscape paintings by utilizing a shuffle attentional residual block and edge enhancement techniques.
For the third type, the output image is guided by textual input information. A novel system Yuan et al. (2023) for Polaca transforms classical poetry into corresponding artistic landscape paintings and calligraphic works. Another approach, controlled landscape painting generation (CCLAP) Wang et al. (2023), is based on the latent diffusion model (LDM) Rombach et al. (2022) and comprises two sequential modules: a content generator and a style aggregator. Although this method allows text-based control of the image, it still faces challenges related to the randomness of image generation and limited controllability.
@@ 2.2 Diffusion-Based Method
DM Ho et al. (2020) is a deep learning framework primarily utilized for image processing and computer vision tasks. Fundamentally, DM work by corrupting the training data by continuously adding Gaussian noise, and then learning to recover the data by reversing this noise process. DM can generate detailed images from textual descriptions and are also effective for image restoration, image drawing, text-to-image, and image-to-image tasks Lin et al. (2023); Zhang et al. (2024a, b); Fei et al. (2023). Essentially, by providing a textual description of the desired image, SD can produce a realistic image that aligns with the given description Ho et al. (2020). These models reframe the "image generation" process into a "diffusion" process that incrementally removes noise. Starting with random Gaussian noise, the models progressively eliminate it through training until it is entirely removed, yielding an image that closely matches the text description Ho et al. (2020); Zhang et al. (2024a). However, a significant drawback of this approach is the considerable time and memory required, especially for high-resolution image generation Lin et al. (2023). LDM Rombach et al. (2022) was developed to address these limitations by significantly reducing memory and computational costs. This is achieved by applying the diffusion process within a lower-dimensional latent space instead of the high-dimensional pixel space Lin et al. (2023).
The emergence of large text-to-image models that can create visually appealing images from brief descriptive prompts has highlighted the remarkable potential of AI. However, these models encounter challenges such as uneven data availability across specific domains compared to the generalized image-to-text domain. In addition, some tasks require more precise control and guidance than simple prompts can provide. ControlNet Zhang et al. (2023) addresses these issues by generating high-quality images based on user-provided cues and controls, which can fine-tune performance for specific tasks. Meanwhile, IP-Adapter Ye et al. (2023) employs decoupled cross-attention mechanisms for the characteristics of the text and the image. Inspired by these approaches, our study incorporates ControlNet’s additional structural controls and the IP-Adapter’s style guiding the generation of landscape paintings.
@@ 3 Proposed Method
To address the above issues, LPGen has been developed to generate landscape paintings from canny image and structure image. This section provides an overview of the framework, the structure controller, which manages the structure of the generated image, and the style controller, which dictates the painting style of the resulting image.
Figure 2: Schematic of LPGen for landscape painting generation The schematic comprises two key components: the structure controller and the style controller. The structure controller module utilizes the Decoupled Cross-Attention technique to separately manage the structural information of an image across different domains, allowing for precise control and regulation of specific elements and attributes in the generated image. The style controller dynamically adjusts the features of the input image, enabling the generative model to accurately capture and reflect the style and structure of the source image.
@@ 3.1 Preliminaries
The complete structure of the LPGen framework is illustrated in Figure 2. The framework comprises three main components: the stable diffusion model, the structure controller, and the style controller.
While the stable diffusion model offers generalized image generation capabilities, it lacks precise control over the structural and stylistic features of an image. To address the structural control problem, we use a structural controller to ensure edge information and effectively guide the image generation process. The structural controller determines the position and shape of graphic elements by outlining, so that the generated image is aligned with the structure of the refined reference image. For stylistic aspects, the style controller extracts the style of a reference segmented image and applies it to the generated image, allowing the generation of images in a specified style. The style controller learns various color schemes and brush features to convey different emotions.
Building upon the stable diffusion model’s basic image generation capabilities, the structure controller manages the structural layout, and the style controller governs the color scheme features. This precise and robust control enables us to achieve excellent management over the style and structural of the generated images, surpassing the capabilities of simple text-to-image models.
@@ 3.2 Structure Controller
The outlining serves as the fundamental framework of landscape painting, establishing the overall layout and positioning of elements. To enable the canny like as outlining of Hand-drawing Method to guide LDM in generating the image, the structure controller manipulates the neural network structure of the diffusion model by incorporating additional conditions. This, in combination with Stable Diffusion, ensures accurate spatial control, effectively addressing the spatial consistency issue. The image generation process thus emulates the actual painting process, transitioning from outlining to coloring. We employ the structure controller as a model capable of adding spatial control to a pre-trained diffusion model beyond basic textual prompts. This controller integrates the UNet architecture from Stable Diffusion with a trainable UNet replica. This replica includes zero convolution layers within the encoder and middle blocks. The complete process executed by the structure controller is as follows:
@@ 3.3 Style Controller
During the coloring stage, the artist sequentially fills and renders the picture, coloring each element according to the guidelines established by the outlines. In the original SD base model with text-to-image generation capability, we introduce the style controller structure to integrate prompt features, structure features, and style features. In order to achieve the above functions, our novel design is a decoupled cross-attention mechanism as shown in Figure 2 , where image features are embedded through a newly added cross-attention layer, which consists mainly of two modules: an image encoder for extracting image features, and an adaptation module with decoupled cross-attention for embedding the image features into a pre-trained text-to-image diffusion model.
The pre-trained CLIP image encoder model is used, but here, in order to efficiently decompose the global image Embedding, we use a small trainable projection network to project the image Embedding into a sequence of features of length N=4. The projection network here is designed as a linear layer Linear plus a normalization layer LN, and at the same time, the dimensions of the input image features are kept consistent with the dimensions of the text features in the pre-trained diffusion model.
@@
$\mathcal{Z}_{new}=\mathcal{\textit{Attention}}(Q,K^{t},V^{t})+\lambda\cdot\mathcal{\textit{Attention}}(Q,K^{i},V^{i}).$ (2)